# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ 3.4 Data Description
The THBS budget book has two primary sections: expenditure and income. I digitized records in each category following the categorization used in the official report of the Tsukishima Survey. Panel A of Table 3 lists the $11$ subcategories for expenditures: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous. These variables are used to test the smoothness of these consumptions. I consider five net expenditure categories: net savings, net insurance, net borrowing, net credit purchase, and net gifts to investigate the risk-coping strategies. Net variables are defined as the expenses minus income in each category; for example, net savings refers to the difference between the deposits to savings minus the withdrawal amount. I also consider the sales of miscellaneous goods and labor earnings from the family members except for the head to test the contribution of the asset sales and intrahousehold labor-supply adjustments. Panel B of Table 3 summarizes these variables. [^47] Online Appendix B.9 presents the results for the panel unit root tests. The null hypothesis of unit roots in all the panels is rejected at the conventional level for all these variables.
The budget book has information on the number of family members. Fortunately, the archives of the Tsukishima Survey include the sheets documenting the changes in family structure (Kazoku idō hyō). [^48] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased). According to the documents, there was only one household experienced changes in the number of family members due to the birth of a son during the sample periods. Therefore, the size of households had been relatively stable over time, suggesting that the preference shifts had rarely occurred in the THBS households. [^49] In village economies, landholding and asset stocks such as livestock and grain may be correlated with the household’s risk-coping behaviors (e.g., Udry 1994; 1995). Since this study analyzes urban factory workers’ households, the roles of these assets can be discarded. Although I include the family size in all the regressions to be conservative, including the family size variable does little to influence the results. This supports the evidence that the preference shifts do not cause endogenous issues in my empirical setting. [^50] Note that the small within variations in the family size variable mean that this control is omitted in some regressions. The number of family members in different age bins is also available from the THBS. Although the shares of family members in each age bin change as time passes (given the birthdays are fixed), these family size variables are close to time-constant. This means that including these variables causes near-perfect multi-collinearity in the fixed-effects models, and thus, it is practically impossible and unnecessary to include these variables.
@@
Total expenditure 34.0 300 3.4 300 61.3 168 4.0 168
    Food 10.8 300 2.2 281 19.4 168 2.8 158
    Housing 2.7 300 1.6 138 4.9 168 1.8 122
    Utilities 1.2 300 -0.1 241 2.1 168 0.4 153
    Furniture 0.4 300 -1.6 187 0.7 168 -1.2 132
    Clothes 2.8 300 0.3 258 5.1 168 1.0 154
    Education 0.7 300 -0.7 212 1.3 168 -0.2 141
    Medical 1.8 300 0.2 280 3.1 168 0.8 158
    Entertainment 1.4 300 -0.1 265 2.5 168 0.5 156
    Transportation 0.4 300 -1.4 203 0.7 168 -0.9 130
    Gift 3.0 300 0.5 252 5.5 168 1.1 154
    Miscellaneous 0.5 300 -1.3 233 0.9 168 -0.9 146
Disposable income 37.4 300 3.5 288 67.1 168 4.1 165
@@
    Net savings 0.23 4.46 300 0.41 6.38 168
    Net insurance -0.10 6.50 300 -0.18 8.75 168
    Net borrowing 0.28 4.93 300 0.56 7.18 168
    Net credit purchase -2.01 11.41 300 -3.84 19.52 168
    Net gifts -1.97 6.01 300 -3.45 9.06 168
@@
    Other family member’s earning 3.48 7.87 300 6.26 14.61 168
    Sales of assets 0.02 0.11 300 0.04 0.15 168
@@
Number of family members 4.3 1.6 2.0 8.0 300
@@
Number of days with illness (head) 0.1 0.8 0 11 263
Number of days with illness 0.5 2.7 0 16 263
@@
Notes: Panel A: The summary statistics of the variables used to test the consumption smoothing for the 34-panel units in the semi-monthly and adjusted monthly panel datasets are shown in this panel. Disposable income is the total income minus tax payments. Panel B: This panel shows the summary statistics of the variables used to test the risk-coping strategies for the 34-panel units in the semi-monthly and adjusted monthly panel datasets. Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit redeemed minus the credit expenses in the credit purchase. “Net gifts” refer to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. Panel C: The summary statistics of the family size variable for the 34-panel units in the semi-monthly panel dataset are listed. The statistics for the adjusted monthly panel dataset are not reported because they are materially similar to those for the semi-monthly panels. The proportions of children in different age bins (%) are not included in the regressions due to multi-collinearities. Panel D: The summary statistics of the number of days with illness for family members are listed. Since this is a lagged variable, a household with two semi-month observations is excluded. Similarly, the initial time period in all the units is not included. Sources: Created by the author from the THBS sample.
@@
Next, I overview the time-series figures to evaluate the representativeness of my THBS dataset. Figure 4c shows an increasing trend in the average income. This is consistent with the macroeconomic trend in the nominal income during this period. This upbeat trend in earnings was, however, offset by the steep inflation under the expansions after the war (Section 2.1). Figure 4d illustrates the variations in income minus expenditure by semi-months. The deviation of expenditure from income in January captures the expenses on the goods for the New Year events. The gaps in June to the first half of July may reflect the payments of bonuses. The humps in March reflect the preparation for the new fiscal and academic year from April, meaning that the increased income in this month may suggest a higher proportion of temporary income from withdrawals and loans. Overall, these support the evidence that the budget data from the THBS households can capture the macroeconomic trends.
@@
Regardless of favorable or adverse shocks, these cells that deviated from mean tendency are helpful in investigating how those households coped with the idiosyncratic shocks. Despite this, it is useful to look into the illness, which is the least predictable shock (Gertler and Gruber 2002). The THBS budget books suggest that the heads had rarely changed their occupation during the survey period. This may be because the target households in this study are the skilled workers’ households who were less likely to move to the other factories. [^52] Another possibility is that the THBS heads who changed their jobs might have been omitted from the survey. However, I confirmed that the households with longer and shorter time-series observations have similar family characteristics (Online Appendix B.3). Thus, this omission does not lead to selection bias in terms of the potential vulnerabilities to the idiosyncratic risk. In contrast, the illness among the THBS family is documented systematically. As explained in Section 2.3, the comprehensive health insurance system had not been available at that time. Moreover, the sanitary conditions in the cities were low, leading to higher mortality rates. [^53] See, for example, Schneider and Ogasawara (2018) for the overviews on the sanitary conditions in interwar Japan. The official report indeed describes that the workers in the machinery factories in Tsukishima had faced the risk of diseases (Department of Health, Ministry of the Interior 1923a, p. 14).
Considering this, I first provide the overall results from the specification using the changes in earnings as the idiosyncratic income shock in Sections 4.1 and 4.2. I then consider the structure behind the adverse income shocks using the heads’ temporary illness to investigate the households’ consumption and risk-coping behaviors. Panel D of Table 3 shows the summary statistics on the health shocks variables used in Section 4.3.
@@
The THBS focuses on the skilled factory workers’ households in the machinery sector in Tsukishima, meaning that this study investigates consumption smoothing behaviors among homogeneous households. This empirical setting is useful because the different attitudes toward risks among households may lead to bias in the standard risk-sharing regressions. Following Schulhofer-Wolf (2011), I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks. Specifically, I consider the heads who worked in large-scale or governmental factories to have potentially different risk preferences because those workers might have lower risk tolerance than those who worked in smaller enterprises. Conditional on the household fixed effects, I confirmed that the sensitivities to the aggregate shocks are materially similar in both groups in the statistical sense. [^54] By using the fixed-effect model with one-way error component, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and government factories, and family size variable. The results are unchanged if I trim the semi-months with the smaller number of cross-sectional observations. This result is unchanged if I assume that a few heads who worked in the smithing factory or in the non-machinery sectors had different risk preferences. [^55] The skilled workers in the smithing sector might be a business owner (Online Appendix A.6). Similarly, the workers in the non-machinery sectors in Tsukishima, where a representative machinery production area, might have had a different, presumably higher, risk-tolerance. This supports the evidence that conditional on the household fixed effects, it is plausible to assume that the THBS households shared similar risk preferences. Online Appendix B.10 summarizes the finer details of this analysis.
@@ 4.1 Consumption Smoothing
I consider a linear fixed-effects model designed to test the consumption smoothing. [^56] Another empirical specification used to test the consumption smoothing is the first-difference model, including the aggregate consumption to capture the macroeconomic shocks in the economy (Mace 1991). In other words, this alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption. Although the consumption of the THBS households may capture the macroeconomic trend in the consumption (Section 3.3), those households do not cover entire households in the target economy, Tsukishima, which shall disturb the interpretation of the coefficient on the aggregate consumption. To be conservative, therefore, I use the two-way fixed-effects model to control for the macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997). Online Appendix C.1 provides the conceptual framework and the derivation of my empirical specification. For household $i$ in time $t$, the reduced form equation is characterized as follows:
@@
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ indicates the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the semi-month fixed effect, and $u_{i,t}$ is a random error term. The household fixed effect captures the time-constant unobservable factors such as permanent income and consumption preference. The time-fixed effect effectively removes the unobservable macroeconomic shocks and trends described in Section 3.4. The estimate of $\gamma$ suggests income elasticity, which can range from zero for perfect insurance to one for the absence of insurance. The cluster-robust variance-covariance matrix estimator is used to deal with heteroskedasticity and serial dependency (Arellano 1987).
@@ Results
Total consumption 0.350 [0.030]*** 288 0.532 [0.071]*** 165 0.271 [0.053]*** 149
    Food 0.130 [0.029]*** 269 0.427 [0.090]*** 155 0.110 [0.068] 140
    Housing 0.132 [0.411] 138 0.311 [0.193] 122 -0.295 [0.159]* 131
    Utilities -0.013 [0.129] 231 0.486 [0.154]*** 150 0.180 [0.302] 138
    Furniture 0.405 [0.219]* 178 0.773 [0.351]** 129 0.325 [0.317] 119
    Clothes 0.432 [0.196]** 247 0.376 [0.326] 151 0.990 [0.304]*** 140
    Education 0.035 [0.089] 204 0.298 [0.194] 139 0.098 [0.201] 130
    Medical expenses 0.074 [0.075] 268 0.507 [0.184]** 155 -0.032 [0.124] 140
    Entertainment expenses 0.299 [0.129]** 253 0.369 [0.180]** 153 -0.072 [0.187] 139
    Transportation 0.288 [0.107]** 195 0.628 [0.183]*** 129 0.371 [0.266] 121
    Gift 0.494 [0.099]*** 242 0.540 [0.131]*** 152 0.595 [0.288]** 138
    Miscellaneous -0.083 [0.154] 224 0.512 [0.175]*** 143 0.231 [0.250] 134
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 1: the regressions of the 11 measures of log-transformed consumption on log-transformed disposable income as well as on the family size control, household-fixed effect, and time-fixed effect. Columns 1, 2, and show the results for the semi-monthly, adjusted monthly and unadjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the column named “Coef.”.
@@
Column (1) of Table 4 presents the results from the semi-monthly panel dataset. In the total consumption, the estimated coefficient is $0.350$. This means that a one percentage point decrease in disposable income decreases the total expenditure by $0.35$ percentage point, suggesting that the factory worker households could not thoroughly smooth their consumptions. Among the subcategories, the food, furniture, clothes, entertainment, transportation, and gifts expenditures are not well-insured. This implies that the households are likely to cut off (increase) their expenses on these categories when they face adverse (favorable) income shocks. The estimates suggest that the expenses on furniture and clothes are relatively sensitive to the shocks ($0.405$ and $0.432$, respectively). This is consistent with the fact that both categories are basically luxury categories. Similarly, the estimate for gifts shows a large magnitude ($0.494$) because the expenses on the other households are essentially optional. The estimates for the entertainment and transportation expenses are slightly smaller but statistically significant ($0.299$ and $0.288$). This also makes sense, given that these expenses are related to family travel and theatergoing. [^57] The workers can enter their factories by walking in Tsukishima. Thus, the costs of commuting to the factory are not usually documented in the budget books.
Importantly, the estimate for food categories is statistically significant but much smaller ($0.130$). This implies that the households might have had some means to deal with idiosyncratic shocks to smooth their expenses on food. I will return to this point later.
@@
Next, I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed in monthly frequency, a longer span than the semi-monthly frequency. Not surprisingly, the adjusted monthly series yields greater estimates than the semi-monthly series. The estimate for total consumption is now $0.532$, which is roughly $1.5$ times greater than that from the semi-monthly panel data. Generally, the panels with short-run time bins are more likely to be influenced by unexpected shocks in consumption, which tends to attenuate the income elasticity (Nelson 1994). In this current case, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels across almost all of the subcategories. The clothes category alone shows a similar but statistically insignificant estimate. This may reflect the cases in which the workers bought their work clothes for the new season. Although such a replacement of work clothes has not occurred frequently (i.e., semi-monthly), this means that the clothes category includes both luxury and necessary items, which may increase the standard errors of the estimate.
@@
A few subcategories turn out to be statistically significantly positive: utilities and medical expenses. The estimate for the utility category is relatively large. Given that the sample period includes the winter season, this reflects the households’ behavior of reducing their expenses on firewood and charcoals in their economic hardship because the firewood and charcoal prices were relatively high. The medical categories include the fee for bathing (yuya), typical for working-class people. This indicates that reducing the frequency of taking baths was relatively easy. By contrast, a few subcategories are not sensitive to the idiosyncratic shocks over different time-series frequencies. The estimates for the housing and education categories are not statistically significant in both semi-monthly and adjusted monthly series. This may reflect the fact that the rent payment and school fees were fixed and paid once per month. [^58] Online Appendix C.2 confirmed this feature in the housing category. In addition, the payments for education shared a very small amount of households’ expenses (Panel A of Table 3).
@@
Overall, I found that the estimated income elasticity is smaller in the semi-monthly than in the adjusted monthly series. In the total consumption, the estimated coefficient is $0.36$ in the semi-monthly panels, while that is $0.56$ in the adjusted monthly panels. One potential explanation behind the part of this difference is that the available sources of temporary income have been limited, and those sources are not useful for compensating the large negative income shocks lasting for at least one month. [^59] The positive income shocks might not be enough to sustain a comfortable lifestyle in the middle run. Given borrowing constraints and a limited amount of liquidity assets, the household could compensate for their income losses for a few weeks but could not do so for several weeks. The most noticeable example is the food category, which comprises the largest portion of the expenditure (Panel A of Table 3). The estimate for the food category is $0.13$, which is more than three times smaller than the estimate from the adjusted monthly series ($0.44$). This implies that the households’ food consumption is likely to be insured in a few weeks but not over one month. I will delve into the mechanism behind this behavior in the following subsections.
@@
Finally, I assess the validity of the aggregation used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (3). The estimate for the total consumption is $0.271$, roughly half of the estimate from the adjusted series ($0.532$). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2) and are statistically insignificant. These estimates are clearly odd, with several subcategories taking values close to one or large negative values, and thus, are no longer be economically interpretable. [^60] For example, the estimate for housing is now statistically significantly negative. The households paid a fixed rent amount monthly. Given that the rent payday is fixed, for example, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these particular months may have created a strong but false positive correlation. This means that measurement errors in the unadjusted aggregation lead to attenuation bias, which results in false negatives in most of the consumption categories and may cause critical issues in statistical inference for a few subcategories. I will return to this point in Section 5 later.

### 25 Feb 2024 v6

@@ 3.4 Data Description
The THBS budget book has two primary sections: expenditure and income. I digitized records in each category following the categorization used in the official report. Panel A of Table 3 lists the $11$ subcategories for expenditures: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous. These are used to test the smoothness of the consumptions in each subcategory. To test the roles of risk-coping strategies, I consider five net income categories: savings, insurance, borrowing, credit purchase, and gifts. Net variables are defined as the income minus expenditure in each category; for example, net savings refers to the difference between the withdrawal amount and deposits to savings. I also consider the sales of miscellaneous goods and labor earnings from the family members except for the head to test the contribution of the asset sales and intrahousehold labor-supply adjustments. Panel B of Table 3 summarizes these variables. [^46] Online Appendix B.9 presents the results for the panel unit root tests. The null hypothesis of unit roots in all the panels is rejected at the conventional level for all these variables.
Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017). [^47] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping behaviors (e.g., Udry 1994, 1995). Since this study analyzes urban factory workers’ households, the roles of these assets can be discarded. Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō). [^48] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased). I found that the size of households had been stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods. This supports the evidence that the preference shifts had rarely occurred among the THBS households. Although family size is included in all regressions to be conservative, it has little effect on the results. [^49] The family size variable shall be omitted in several regressions due to the near-perfect multi-collinearity in the fixed-effects models given its small within variations. Although the number of family members in different age bins is available from the THBS, those family composition variables are also time-constant and are absorbed by the household fixed effect. Panel C of Table 3 shows the summary statistics on the family size variable.
@@
Total expenditure 34.5 289 3.4 289 61.1 162 4.0 162
    Food 10.9 289 2.3 270 19.5 162 2.8 152
    Housing 2.7 289 1.6 134 4.9 162 1.8 118
    Utilities 1.2 289 -0.1 232 2.2 162 0.5 147
    Furniture 0.4 289 -1.6 177 0.7 162 -1.2 126
    Clothes 2.8 289 0.3 248 5.0 162 1.0 148
    Education 0.7 289 -0.6 201 1.3 162 -0.2 135
    Medical 1.7 289 0.2 269 3.0 162 0.7 152
    Entertainment 1.4 289 -0.1 255 2.5 162 0.5 150
    Transportation 0.4 289 -1.3 195 0.8 162 -0.9 125
    Gift 3.1 289 0.5 243 5.6 162 1.2 148
    Miscellaneous 0.5 289 -1.3 222 0.9 162 -0.9 140
Disposable income 37.6 289 3.5 278 67.3 162 4.1 159
@@
    Net savings -0.01 3.97 289 -0.02 5.52 162
    Net insurance -0.03 6.67 289 -0.05 8.97 162
    Net borrowing 0.31 5.03 289 0.61 7.32 162
    Net credit purchase -2.47 11.33 289 -4.67 19.38 162
    Net gifts -2.04 6.11 289 -3.56 9.20 162
@@
    Other family member’s earning 3.74 6.11 289 6.72 14.79 162
    Sales of assets 0.02 0.12 289 0.04 0.15 162
@@
Number of family members 4.3 1.6 2.0 8.0 289
@@
Number of days with illness (head) 0.1 0.8 0 11 254
Number of days with illness 0.2 1.7 0 16 254
@@
Notes: Panel A: The summary statistics of the variables used to test the consumption smoothing for the 33-panel units in the semi-monthly and adjusted monthly panel datasets are shown in this panel. Disposable income is the total income minus tax payments. Panel B: This panel shows the summary statistics of the variables used to test the risk-coping strategies for the 33-panel units in the semi-monthly and adjusted monthly panel datasets. Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. Panel C: The summary statistics of the family size variable for the 33-panel units in the semi-monthly panel dataset are listed. The statistics for the adjusted monthly panel dataset are not reported because they are materially similar to those for the semi-monthly panels. Panel D: The summary statistics of the number of days with illness for family members are listed. Since both are lagged variables, the initial time period in all the units is omitted. Similarly, a household with two semi-month observations is excluded. Similarly, . Sources: Created by the author from the THBS sample.
@@
Next, I illustrate the time-series figures of the THBS dataset to overview the trend in income and expenditure. Figure 4c shows a weak increasing trend in the average income, which is consistent with the overall macroeconomic trend in the nominal income during this period. This trend in earnings was, however, offset by the steep inflation under the expansions after the war (Section 2.1). In Figure 4d, the deviation of expenditure from income in January must capture the expenses on the goods for the New Year events. The gaps in June to the first half of July may reflect the payments of bonuses. The humps in March may reflect the preparation for the new fiscal and academic year from April, suggesting a higher proportion of temporary income from withdrawals and loans. These support the evidence that the budget data from the THBS households can capture a large part of the macroeconomic trends.
@@
Considering this, I first provide the overall results from the specification using the changes in earnings as the idiosyncratic income shock in Sections 4.1 and 4.2. I then consider the structure behind the adverse income shocks using the heads’ temporary illness to investigate the households’ consumption and risk-coping behaviors in Section 4.3. Panel D of Table 3 shows the summary statistics on the health shock variables.
@@
This study investigates consumption smoothing behaviors among the skilled factory workers’ households in the machinery sector in Tsukishima. This empirical setting is preferable because the different attitudes toward risks among households with different occupations and regions may lead to bias in the standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). To assess the homogeneity of the risk preference among THBS households, I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks in the spirit of Schulhofer-Wohl (2011). Specifically, I consider the heads who worked in large-scale or governmental factories to have potentially different risk preferences because those workers might have lower risk tolerance than those who worked in smaller enterprises. I confirmed that the sensitivities to the aggregate shocks are materially similar in both groups in the statistical sense. [^53] Online Appendix B.10 summarizes the finer details of this analysis. In short, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and government factories, family size variable, and household fixed effects. The results are unchanged if I trim a few semi-months with the smaller number of cross-sectional observations. This result is unchanged if I assume that the heads who worked in the smithing factory or in the non-machinery sectors had different risk preferences. [^54] The skilled workers in the smithing sector might be a business owner (Online Appendix A.6). Similarly, the workers in the non-machinery sectors in Tsukishima, a representative machinery production area, might have had a different, presumably higher, risk-tolerance. This supports the evidence that conditional on the household fixed effects, it is plausible to assume that the THBS households shared similar risk preferences.
@@ 4.1 Consumption Smoothing
I consider a linear fixed-effects model designed to test the consumption smoothing. [^55] Another empirical specification used to test the consumption smoothing is the first-difference model, including the aggregate consumption to capture the macroeconomic shocks in the economy (Mace 1991). In other words, this alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption. Although the consumption of the THBS households tends to capture the macroeconomic trend in the consumption (Section 3.3), those households do not cover entire households in the target economy, Tsukishima. This disturbs the interpretation of the coefficient on aggregate consumption. To be conservative, therefore, I use the two-way fixed-effects model to control for the macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997). Online Appendix C.1 provides the conceptual framework and the derivation of my empirical specification. For household $i$ in time $t$, the reduced form equation is characterized as follows:
@@
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the semi-month fixed effect, and $u_{i,t}$ is a random error term. The household fixed effect captures the time-constant unobservable factors such as permanent income and consumption preference. The time-fixed effect effectively removes the unobservable macroeconomic shocks and trends described in Section 3.4. The estimate of $\gamma$ shows income elasticity, which can range from zero for perfect insurance to one for the absence of insurance. The cluster-robust variance-covariance matrix estimator is used to deal with heteroskedasticity and serial dependency (Arellano 1987).
@@ Results
Total consumption 0.358 [0.033]*** 278 0.548 [0.065]*** 159 0.261 [0.054]*** 144
    Food 0.136 [0.030]*** 259 0.436 [0.088]*** 149 0.100 [0.072] 135
    Housing 0.099 [0.415] 134 0.291 [0.243] 118 -0.263 [0.167] 126
    Utilities -0.003 [0.138] 223 0.431 [0.149]*** 144 0.197 [0.333] 133
    Furniture 0.426 [0.223]* 169 0.943 [0.335]*** 123 0.250 [0.339] 114
    Clothes 0.378 [0.190]* 238 0.374 [0.324] 145 0.806 [0.262]*** 135
    Education 0.070 [0.092] 194 0.337 [0.212] 133 0.093 [0.213] 125
    Medical expenses 0.039 [0.066] 258 0.498 [0.196]** 149 -0.128 [0.094] 135
    Entertainment expenses 0.355 [0.109]*** 244 0.484 [0.148]*** 147 -0.049 [0.199] 134
    Transportation 0.278 [0.110]** 187 0.635 [0.190]*** 124 0.362 [0.290] 117
    Gift 0.523 [0.106]*** 234 0.599 [0.115]*** 146 0.622 [0.305]* 133
    Miscellaneous -0.076 [0.147] 214 0.549 [0.184]*** 137 0.246 [0.267] 129
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 1: the regressions of the 11 measures of log-transformed consumption on log-transformed disposable income as well as on the family size control, household-fixed effect, and time-fixed effect. Columns 1, 2, and 3 show the results for the semi-monthly, adjusted monthly, and unadjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
@@
Column (1) of Table 4 presents the results from the semi-monthly panel dataset. In the total consumption, the estimated coefficient is $0.358$. This means that a one percentage point decrease in disposable income decreases the total expenditure by $0.36$ percentage point, suggesting that the factory worker households could not thoroughly smooth their consumptions. Among the subcategories, the food, furniture, clothes, entertainment, transportation, and gifts expenditures are not well-insured. The households are likely to cut off (increase) their expenses on these categories when they face adverse (favorable) income shocks. The estimates suggest that the expenses on furniture and clothes are relatively sensitive to the shocks ($0.426$ and $0.378$, respectively), which is consistent with the fact that both categories are luxury categories. Similarly, the estimate for gifts shows a large magnitude ($0.523$), reflecting that the expenses on the others are essentially optional. The estimates for the entertainment and transportation expenses are slightly smaller but statistically significant ($0.355$ and $0.278$). This also makes sense, given that these expenses include family travel and theatergoing. [^56] The workers can enter their factories by walking in Tsukishima. Thus, the costs of commuting to the factory are not usually documented in the budget books. An important finding is that the estimate for food categories is statistically significant but much smaller ($0.136$). Therefore, the households might have had some means to deal with idiosyncratic shocks to smooth their expenses on food. I will return to this point later.
@@
I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed in a longer span. Not surprisingly, the adjusted monthly series yields greater estimates than the semi-monthly series. The estimate for total consumption is now $0.548$, which is roughly $1.5$ times greater than that from the semi-monthly series. Generally, the panels with short-run time bins are more likely to be influenced by unexpected shocks in consumption, which tends to attenuate the income elasticity (Nelson 1994). If this mechanism worked, subcategories that are sensitive (insensitive) to unexpected shocks should be more (less) attenuated. In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels across almost all subcategories. Thus, this suggests that households may face more difficulties in smoothing their consumption in the longer run. The clothes category alone shows a similar but statistically insignificant estimate. This may reflect the cases in which the workers bought their work clothes for the new season. Since work clothes are rather necessary items, this sort of seasonal expense may attenuate the estimate.
@@
A few subcategories turn out to be statistically significantly positive: utilities and medical expenses. The estimate for the utility category is relatively large. Given that the sample period includes the winter season and that the firewood, charcoals, and medicines are relatively expensive, this may reflect the households’ behavior of reducing their expenses on fuels and medicines. By contrast, a few subcategories are not sensitive to the idiosyncratic shocks over different time-series frequencies. The estimates for the housing and education categories are not statistically significant in both semi-monthly and adjusted monthly series. This may reflect that the rent payment and school fees were fixed and paid once per month. [^57] Online Appendix C.2 confirmed this feature in the housing category. In addition, the payments for education shared a very small amount of households’ expenses (Panel A of Table 3).
@@
Overall, I found that the estimated income elasticity is smaller in the semi-monthly than in the adjusted monthly series. In the total consumption, the estimated coefficient is $0.36$ in the semi-monthly panels, while that is $0.55$ in the adjusted monthly panels. One potential explanation behind this difference is that the available sources of temporary income have been limited, and those sources are not useful for compensating for adverse income shocks lasting for at least one month. [^58] The positive income shocks might not be enough to sustain a comfortable lifestyle in the middle run. In other words, while the household could compensate for their income losses for a few weeks, they could not do so for several weeks, presumably because of borrowing constraints and/or a limited amount of liquidity assets. The most noticeable example is the food category, which comprises the largest portion of the expenditure (Panel A of Table 3). The estimate for the food category is $0.14$, which is more than three times smaller than the estimate from the adjusted monthly series ($0.44$). In addition, the estimates for the utilities and medical expenses show a relatively clear tendency. This implies that the households’ expenditures on food, fuel, and medicine are likely to be insured in a few weeks but not over one month. As I will delve into the mechanism behind this behavior in the following subsections, the credit purchases may work to smooth consumptions for these categories in the short run.
@@
Finally, I assess the validity of the aggregation method used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (3). The estimate for the total consumption is $0.261$, roughly half of the estimate from the adjusted series ($0.548$). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2). These estimates are clearly odd, with several subcategories taking values close to one or large negative values. Those are no longer economically interpretable. [^59] For example, the estimate for housing is now negative. The households paid a fixed rent amount monthly. Given that the rent payday is fixed, for example, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these particular months may have created a strong but false positive correlation. This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives in most of the consumption categories and may cause critical issues in statistical inference for a few subcategories.